Reports 1-1 of 1 Clear search Modify search
AdV-TCS (CO2 laser projector)
mantovani, nardecchia, pinto, ruggi, spinicelli - 17:48 Friday 22 September 2023 (61754) Print this report
DAS tuning

After solving the problem on the BS DSP, DAS tuning started.

9.10 UTC: LN3

9.35 UTC: WI DAS INCREASE (+15 %)

10.01UTC: unlock

10.53 UTC: the sidebands was low in the CITF preventing the lock acquisition→ increase WI CH by 15 %

The powers applied in this step are summarized in the table below.

 

STEP6_S

12.39 UTC

CH [W]

INNER DAS [W]

OUTER DAS [W]

W

on the ITF

0.078

0.090*(1+15/100)=0.119

1.615*(1+15/100)=2.45

on the pickoff

0.48

0.0169*(1+15/100)= 0.0194

0.347*(1+15/100)=0.399

N

on the ITF

0.091

0.266 

2.972

on the pickoff

0.561

0.044

0.484

After that, it was possible to reach CARM NULL 1F but the misbalance of the upper and lower 56 MHz sidebands on B1p prevented the lock in low noise, thus I reverted the increase of the WI DAS power and, following Paolo’s indication, I slightly decreased the NI DAS by 10 % (at 13.28 UTC).

 

STEP7_S
13.28 UTC

CH [W]

INNER DAS [W]

OUTER DAS [W]

W

on the ITF

0.078

0.104

2.13

on the pickoff

0.48

0.0169

0.347

N

on the ITF

0.091

0.186*(1-10/100)+0.08=

0.247 

1.172*(1-10/100)+1.8 =2.85

on the pickoff

0.561

0.036*(1-10/100)+0.013=

0.0402

0.190*(1-10/100)+0.293=

0.464

 

The ITF unlocked during different steps of the lock acquisition, thus, the DAS powers have been re-settled to STEP 5S (at 15.22 UTC).

 

STEP5_S 

CH [W]

INNER DAS [W]

OUTER DAS [W]

W

on the ITF

0.068

0.104

2.13

on the pickoff

0.418

0.0169

0.347

N

on the ITF

0.091

0.266 

2.972

on the pickoff

0.561

0.044

0.484

Sorry for the delay of the start of the afteroon activity. 

Comments to this report:
nardecchia, pinto, gouaty, magazzu, tringali - 22:51 Friday 22 September 2023 (61755) Print this report

Today afternoon, it was not possible to reach LN2 most probably due to the unbalance of the B1p 56 MHz upper and lower sidebands (see fig.1).

In the lock started at 18.11 UTC (when CARM NULL 1F has been achieved), we added -11 as MICH offset in order to overimpose the sidebands  to proceed up to LN2, the trick worked (see fig.2) but the ITF unlocked at LN2 due to huge oscillation at ~55 Hz .  

Thus, to save the weekend, we decided to reverse the last step of DAS WP.

So, at 19.24 UTC, we came back to STEP_1S (61729) .

 

STEP_1S
19.24 UTC

CH [W]

INNER DAS [W]

OUTER DAS [W]

W

on the ITF

0.068

0.090

1.857

on the pickoff

0.418

 0.0146

0.302

N

on the ITF

0.091

0.266 

2.972

on the pickoff

0.561

0.044

0.484

After that, the ITF achieved LN3 (at 20.25 UTC) without adding extra offsets (fig.3).

As requested by Romain, the BS TY SET was put to 0.

We left the ITF to Romain.

Images attached to this comment
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×